If 0 log(2xy) = logx logy log(yz) = logy logz log(2zx) = logz logx Find 2x+3y+4z

If 0
log(2xy) = logx logy
log(yz) = logy logz
log(2zx) = logz logx
Find 2x+3y+4z

Grade:

1 Answers

Jayant Kishore
8 years ago
This is only possible when
log(2xy) =log(yz)=log(2zx)=logz logx = logy logz = logx logy=0
Therefore
x=1/2   y=1   z=1 
So,     
2x+3y+4z
= 2X1/2 +3X1+4X1
=8